Grade 12th passAlgebra

The normal to the rectangular hyperbola xy = 8 at the point (4, 2) meets the asymp-
totes at M and N. Find the length of MN.

nartey eugene
7 Years agoGrade 12th pass

1 Answer

Samyak Jain
7 Years ago
Equation of assymptotes to the rectangular hyperbola xy = 8 are x = 0 & y = 0.
Differentiate xy = 8 wrt x  \Rightarrow  x dy/dx + y.1 = 0.      dy/dx = –y/x
 \therefore Slope of tangent at (4,2) is –2/4 = –1/2
\therefore  Slope of normal at (4.2) is –1/(–1/2) = 2.
Equation of normal is y – 2 = 2(x – 4)  \Rightarrow  2x – y – 6 = 0.
It intersects x-axis at M(3,0) and y-axis at N(0,–6).
By distance formula, MN = \sqrt{(3-0)^2 + (0+6)^2}  =  \sqrt{9+36} = \sqrt{45} = 3\sqrt{5}